OpenStudy (anonymous):

sketch the graph of y=2x^2-12x+5 by completing the square and making appropriate transformations of y=x^2

OpenStudy (mathmate):

y=2(x-3)^2-13 So the graph has its vertex at (3,-13), and y-intercept +5. The zeroes are at 3+/-sqrt(6.5)

OpenStudy (anonymous):

can you explain how you did that?

OpenStudy (mathmate):

y=2(x-3)^2-13 is obtained by completing the squares. The y-intercept is from the constant term of the original expression. The zeroes are from solving 2(x-3)^2-13=0

OpenStudy (anonymous):

The sketch of the graph should look something like the attachment.

OpenStudy (mathmate):

Oh, I forgot to mention, the graph is concave upwards, which is obtained by the positive value of +2 of the leading coefficient (2x^2+...) Hi Rob: Maybe it's computer glitch, but I don't see much in the attachment.

OpenStudy (anonymous):

Very strange. The attachment is a Mathematica generated plot of 2x^2-12x+5 between x=0 and x=6 which is definitive ie: Far beyond a human generated sketch. Adjust your PDF viewer so that the you can see the entire curve showing the vertex is at about (3,-13) and the y intercept at (0,5).

OpenStudy (mathmate):

Thanks for you info. For some reason it displays now. Probably my old workhorse took a coffee break. It IS a very nice plot.

OpenStudy (anonymous):

I'm glad that you are convinced. Have a good day.

OpenStudy (mathmate):

Same to you too!